A freight train and an express train leave towns 390 km apart, traveling toward one another. The freight train travels 30 km per hour slower than the express train. They pass one another 3 hours later. What are their speeds? I think the speed of the freight may be 100km/hr and express would be 130km/hr, using the velocity = distance/time equation, but I am not quite sure of the formula and expression, or the answer for that matter. So if you could please help me I would appreciate it.
Thanks
: A freight train and an express train leave towns 390 km apart, traveling toward one another. The freight train travels 30 km per hour slower than the express train. They pass one another 3 hours later. What are their speeds? I think the speed of the freight may be 100km/hr and express would be 130km/hr, using the velocity = distance/time equation, but I am not quite sure of the formula and expression, or the answer for that matter. So if you could please help me I would appreciate it.
Thanks

Answer by ankor@dixie-net.com(4532)

A freight train and an express train leave towns 390 km apart, traveling toward one another. The freight train travels 30 km per hour slower than the express train. They pass one another 3 hours later. What are their speeds?
:
One thing we do know is that the total distance traveled by the two trains = 390km
:
Let s = speed of the freight
then
(s+30) = speed of the express
:
Write a distance equation: Dist = time * speed
:
Freight dist + Express dist = 390 km
3s + 3(s+30) = 390
;
3s + 3s + 90 = 390
:
6s = 390 - 90
:
6s = 300
s = 300/6
s = 50 km/hr speed of the freight
and
50 + 30 = 80 km/hr speed of the express
:
:
Check solution by find the total dist
3(50) + 3(80) = 390
: